Neverbreak, Inc., produces boxes to be used in long-distance shipments. The boxes come in sizes “small” (S) and “large” (L), and in strengths “regular” (R) and “neverbreak” (N). So the company produces four types of boxes: small regular (SR), small neverbreak (SN), large regular (LR), and large neverbreak (LN). The demand for the boxes in next month is forecasted by management as 22 for SR, 40 for SN, 14 for LR, and 5 for LN.

The company has two types of machines. The first type of machine can produce 100 small boxes of any kind per machine each month and requires two operators. There is one of this type of machine available. The second type of machine can produce 30 large boxes of any kind  per machine per month and requires 3 operators to work on it. There are 4 machines of this type available.

What is the number of machines required to produce small boxes next month (Keep 2 decimal points!) ?
Answer #1
  • Total no of small boxes = no. of small regular + no. of small never break = 22+40 = 64
  • Capacity of the machine = 100 boxes
  • No of machines for small boxes = total boxes / capacity of each machine = 64/100= 0.80
  • No of machines required for large boxes = 0.64 (Answer)

****Please please please LIKE THIS ANSWER, so that I can get a small benefit, Please****